# Exact and Linear Convergence for Federated Learning under Arbitrary Client Participation is Attainable

### v2

@@ 4.2.1 Push-Pull strategy for FL setting
Note a minor but critical difference from the formulation (1) is that there are $N+1$ model parameters $x_{i}$ applied in each local cost function $f_{i}$ instead of a single $x$. To see the equivalency between this and (1), notice $R(S_{r}){\boldsymbol{x}}={\boldsymbol{x}}$ implies $x_{i}=x_{0},\forall i\in S_{r}$. Consequently, if the union of all sampled client sets $\{S_{r}\}$ covers the entire client population, then all individual client models and the server model are constrained to converge to the same state.
@@
Algorithm 1
FOCUS
: Federated Optimization with Exact Convergence via Push-pull Strategy
1:  Notation: $x_{r}$ and $y_{r}$ are the server’s values while $x^{(r)}_{t,i}$ and $y^{(r)}_{t,i}$ are client’s values.
@@
6:        $x^{(r)}_{0,i}\Leftarrow x_{r},\;y^{(r)}_{0,i}\Leftarrow 0$triangleright Pull $x_{r}$ (No need to pull $y_{r}$)
@@
8:           $y^{(r)}_{t+1,i}=y^{(r)}_{t,i}+\nabla f_{i}(x^{(r)}_{t,i})-\nabla f_{i}(x^{(r)}_{t-1,i})$triangleright See $\nabla f_{i}(x^{(r)}_{-1,i})$ around (22).
@@
12:     $y_{r+1}\Leftarrow y_{r}+\sum_{i\in S_{r}}y^{(r)}_{\tau,i}$triangleright Push $y_{\tau,i}^{(r)}$ (Not Average)
@@ 4.2.2 Convert Vector-Matrix form back to FL-style Algorithm
$y_{k+1,0}$ $=y_{k,0}+\sum_{i\in S_{r}}y^{\prime}_{k+1,i}$ (server collects info) (18)
@@ 5 Performance Analysis
Now we are ready to present the necessary assumptions and convergence property for FOCUS . Due to limited space, all proofs are deferred to Appendix D.
@@
where the Lyapunov functions $\Psi_{r}:=\mathbb{E}\hskip 1.42271pt\|\bar{x}_{r\tau+1}-x^{\star}\|^{2}+(1-8\eta\tau LN)\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{(r-1)\tau+1}-{\boldsymbol{x}}_{r\tau}\|^{2}_{F}$, $\Phi_{r}=\mathbb{E}\hskip 1.42271ptF(\bar{x}_{r\tau+1})-F^{\star}+\left(1-4\eta L^{2}\right)\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{r\tau+1}-{\boldsymbol{x}}_{r\tau}\|^{2}$ and $q_{\min}=\min_{i}q_{i}$. $\hfill\qed$
@@ 5.2 Extension to Stochastic Gradients and ML Applications
In practical machine learning scenarios, computing full gradients is often computationally prohibitive. Therefore, stochastic gradient methods are commonly employed. Our proposed algorithm can be readily extended to incorporate stochastic gradients, resulting in the variant SG-FOCUS . However, due to space constraints, we focus on the deterministic setting in the main body of this paper. A comprehensive description of SG-FOCUS , along with its convergence analysis, is provided in Appendix E. The appendix also benchmarks SG-FOCUS ’s performance on the CIFAR-10 classification task, highlighting its faster convergence and improved accuracy over other FL algorithms. This performance trend echoes that of its deterministic counterpart.
Figure 4: Convergence performance comparison of various FL algorithms. Under full client participation, FedAvg and FedAU exhibit identical performance, as do SCAFFOLD and ProxSkip, due to their theoretical equivalence in this setting. FedAvg and FedAU fail to converge to the optimal solution across all scenarios because their inherent error and bias cannot be eliminated using a fixed learning rate. ProxSkip diverges under uniform and arbitrary participation, as it is not designed for these conditions. While SCAFFOLD converges in all cases, our proposed algorithm, FOCUS , demonstrates faster convergence, especially under arbitrary participation.
5.3 Limitations and Future works
@@
The arbitrary client participation modeling used in the proof of this paper didn’t consider the Markov process, i.e., the client participation probabilities depend on the participation status in the last round. We believe FOCUS still converges exactly under this scenario since the stochastic matrix modeling and push-pull strategy still hold for any realizations. However, the extension of the proof is non-trivial due to the correlation between stochastic matrices. We leave this and a more general arbitrary participation scenario for future research directions.
@@ Appendix A Review of Decentralized Algorithms
It is well-known that the DGD with a fixed step size $\eta$ only converges to an ${\mathcal{O}}(\eta)$-sized neighborhood of the solution of the original algorithm (Yuan et al., 2016). This resembles the FedAvg algorithm exactly. Subsequently, several extract decentralized optimization algorithms have been proposed, such as Extra (Shi et al., 2015), exact-diffusion/NIDS (Yuan et al., 2018; Li et al., 2019b), DIGing/Gradient tracking (Nedic et al., 2017), etc. The key advancement of these algorithms is their capability to achieve extra convergence under a fixed step size. They formulate the original sum-of-cost problem into a constrained optimization problem and then apply the primal-dual style approach to solve the constrained problem (Ryu and Yin, 2022).
While it is common in the analysis of decentralized algorithms to assume a static and strongly connected underlying graph structure, a significant body of research also investigates time-varying graph topologies (Lan et al., 2020; Saadatniaki et al., 2020; Assran et al., 2019; Ying et al., 2021; Nguyen et al., 2025). These studies often adopt one of two common assumptions regarding the dynamics of such graphs: either the union of graphs over any consecutive $\tau-$ iterations or the expected graph is strongly connected (Nedić and Olshevsky, 2014; Koloskova et al., 2020). Whereas previous research on time-varying topologies often relies on specific graph assumptions, this paper takes a different approach. We model client sampling and local updates using a graph representation, thereby avoiding any presuppositions about the underlying graph structure.
Lastly, we want to point out that this work focuses quite differently from the decentralized FL work Beltrán et al. (2023); Shi et al. (2023); Fang et al. (2024), which is more closely related to decentralized algorithms instead of FL settings.
@@ Appendix B Conventions and Notations
[…]
@@
Appendix C Proof of the Convergence of FedAvg under Arbitrary Client Participation
@@
This section presents a convergence analysis of the FedAvg algorithm with an arbitrary sampling/participation scheme. We focus on the strongly-convex case with a constant step size since this setting best illustrates the impact of client drift induced by local updates and bias introduced by non-uniform sampling. The following proof draws inspiration from and synthesizes several existing works (Wang and Ji, 2022; Koloskova et al., 2020; Li et al., 2019a), adapting their insights to a decentralized framework. Leveraging this framework, we are able to present a more concise proof and provide a clearer conclusion. Unlike FOCUS , most FL algorithms require an extra bounded heterogeneity assumptions:
@@
where $\bar{x}_{k}=q^{\mathsf{T}}{\boldsymbol{x}}_{k}$, $x^{\star}$ is the optimal point of (1) and $\kappa=L/\mu$ is the condition number. $\hfill\qed$
@@
C.1 Reformation and Mixing Matrices
@@
See the definition of $R(S_{r})$ and $V(S_{r})$ in main context and example in Figure 6. A noteworthy observation from (Li et al., 2019a) is that FedAvg can be equivalently reformulated without altering the trajectory of the server’s model. This reformulation considers activating all devices, where each device pulls the model from the server and performs a local update, while maintaining the same set of contributing clients for averaging. See Figure 7 as an illustration. Swap the order of $x$ and $y$ update, we arrive at

### v3

@@ 4.2.1 Push-Pull Strategy for FL Settings
Note a minor but critical difference from the formulation (1) is that there are $N+1$ model parameters $x_{i}$ applied in each local cost function $f_{i}$ instead of a single $x$. To see the equivalence between this and (1), notice $R(S_{r}){\boldsymbol{x}}={\boldsymbol{x}}$ implies $x_{i}=x_{0},\forall i\in S_{r}$. Consequently, if the union of all sampled client sets $\{S_{r}\}$ covers the entire client population, then all individual client models and the server model are constrained to converge to the same state.
@@
Algorithm 1 FOCUS: Federated Optimization with Exact Convergence via Push-pull Strategy  [^3] The code is available at https://github.com/BichengYing/FedASL. The algorithm was originally named Federated Learning for Arbitrary Sampling and Local Update (FedASL). The acronym ASL also stands for the Adaptation System Laboratory at UCLA and EPFL, where Dr. Ying completed his Ph.D.
1:Notation: $x_{r}$ and $y_{r}$ are the server’s values while $x^{(r)}_{t,i}$ and $y^{(r)}_{t,i}$ are clients’ values.
@@
6:   $x^{(r)}_{0,i}\Leftarrow x_{r},\;y^{(r)}_{0,i}\Leftarrow 0$ $\triangleright$ Pull $x_{r}$ (No need to pull $y_{r}$)
@@
8:    $y^{(r)}_{t+1,i}=y^{(r)}_{t,i}+\nabla f_{i}(x^{(r)}_{t,i})-\nabla f_{i}(x^{(r)}_{t-1,i}) \triangleright$ See $\nabla f_{i}(x^{(r)}_{-1,i})$ around (22).
@@
12:  $y_{r+1}\Leftarrow y_{r}+\sum_{i\in S_{r}}y^{(r)}_{\tau,i} \triangleright$ Push $y_{\tau,i}^{(r)}$ (Not Averaged)
@@ 4.2.2 Convert Vector-Matrix Form Back To FL-Style Algorithm
$y_{k+1,0}$ $\Leftarrow y_{k,0}+\sum_{i\in S_{r}}y^{\prime}_{k+1,i}$ (server collects info) (18)
@@ 5 Performance Analysis
Now we are ready to present the necessary assumptions and convergence property for FOCUS. Due to limited space, all proofs are deferred to Appendix E.
@@
[…]
@@
Figure 4: Convergence performance comparison of various FL algorithms. Under full client participation, FedAvg, FedAU, and MIFA exhibit identical performance, as do SCAFFOLD and ProxSkip, due to their theoretical equivalence in this setting. FedAvg and FedAU fail to converge to the optimal solution across all scenarios because their inherent error and bias cannot be eliminated using a fixed learning rate. ProxSkip diverges under uniform and arbitrary participation, as it is not designed for these conditions. We do not understand why MIFA diverges but it works in ML applications. While SCAFFOLD converges in all cases, our proposed algorithm, FOCUS, demonstrates faster convergence, especially under arbitrary participation.
@@ 5.2 Extension to Stochastic Gradients and ML Applications
In practical machine learning scenarios, computing full gradients is often computationally prohibitive. Therefore, stochastic gradient methods are commonly employed. Our proposed algorithm can be readily extended to incorporate stochastic gradients, resulting in the variant SG-FOCUS . However, due to space constraints, we focus on the deterministic setting in the main body of this paper. A comprehensive description of SG-FOCUS , along with its convergence analysis, is provided in Appendix F. The appendix also benchmarks SG-FOCUS ’s performance on the CIFAR-10 classification task, highlighting its faster convergence and improved accuracy over other FL algorithms. This performance trend echoes that of its deterministic counterpart.
Theorem 2 (Informal Convergence Theorem of SG-FOCUS ).
Under arbitrary participation assumption 1, $L$-Smoothness assumption 2, and unbiased and bounded variance assumption on stochastic gradient (See assumption 6 in appendix F), it can be proved that SG-FOCUS converges at the following rates with various extra assumptions on $f_{i}$:
  • $\mu$-Strongly Convex: Under extra assumption 3, for sufficiently small learning rate $\eta$, we have
    $\Gamma_{R}\leq$ $\left(1-\frac{\eta\mu N}{2}\right)^{R}\Gamma_{0}+\frac{4(q_{\min}+N^{2})}{\mu Nq_{\min}}\eta\sigma^{2},$ (26)
    where the Lyapunov functions $\Gamma_{r}:=\mathbb{E}\hskip 1.42271pt\|\bar{x}_{(r+1)\tau+1}-x^{\star}\|^{2}+\big(1-8\eta L^{2}N/\mu\big)\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{(r-1)\tau+1}-{\boldsymbol{x}}_{(r-1)\tau}\|^{2}_{F}$ and $\sigma$ is the variance upper bound of the stochastic gradient noise.
  • $\beta$-PL Condition: Under extra assumption 4, for sufficient small learning rate $\eta$, we have
    $\Omega_{R}\leq$ $(1-\eta\beta N)^{R}\Omega_{0}+\left(\frac{L}{2}+32(\tau-1)^{2}L^{2}+\frac{8}{q_{\min}}\right)\frac{\eta}{\beta}\sigma^{2},$ (27)
    where $\Omega_{r}:=F(\bar{x}_{r\tau+1})-F^{\star}+\left(1-4\eta L^{2}\right)\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{(r-1)\tau+1}-{\boldsymbol{x}}_{(r-1)\tau}\|^{2}$.
  • General Nonconvex: Under no extra assumption, for sufficient small learning rate $\eta$, we have
    $\frac{1}{R}\sum_{r=0}^{R-1}\mathbb{E}\hskip 1.42271pt\|\nabla F(\bar{x}_{r\tau+1})\|^{2}\leq$ $\frac{8\big(F(\bar{x}_{1})-F^{\star}\big)}{\eta NR}+\left(2LN+\frac{8N^{2}}{(N-8L^{2})q_{\min}}\right)\eta^{2}\sigma^{2}.$ (28)
[…]
Remarks on Linear Speedup. A common expectation in FL theory is the demonstration of a linear speedup in the convergence rate, where the rate scales proportionally with the number of clients, $N$. By inspecting the general non-convex convergence rate in Theorem 1, we see the error residual term is ${\mathcal{O}}(1/(\eta NR))$. Yet, the stability of FOCUS necessitates a learning rate $\eta$ that is restricted by ${\mathcal{O}}(1/N)$. This $N$ dependence cancels out. We want to highlight that this result is expected. The linear speedup typically holds when N is used to average out stochastic noise (like in Stochastic Gradient Descent variants). Since FOCUS is an exact algorithm, it does not introduce this stochasticity, and therefore, it is natural that the linear speedup benefit from increasing the number of clients is not reflected in the convergence bound. In contrast to the analysis of FOCUS, the convergence rate of SG-FOCUS in the general non-convex setting does indeed reflect the benefits of client aggregation. Specifically, by setting the learning rate $\eta$ to ${\mathcal{O}}(1/N)$, the stochastic variance term diminishes proportionally to ${\mathcal{O}}(1/N)$, which confirms the presence of the linear speedup.
6 Conclusion
This work addresses the critical challenges of arbitrary client participation and client drift in Federated Learning, two factors that prevent traditional algorithms from achieving exact convergence. By introducing a novel framework based on stochastic matrices and time-varying graphs, we model these dynamics and reformulate the FL problem as a constrained optimization task. This principled approach, moving beyond simple heuristics, led to the development of FOCUS, an algorithm derived from the push-pull optimization strategy. Our theoretical analysis and numerical experiments demonstrate that FOCUS can achieve exact linear convergence under any client participation scheme, without needing to know or learn participation probabilities. The extension to a stochastic gradient setting, SG-FOCUS , further validates its practical effectiveness.
@@ 6 Conclusion
Limitations The arbitrary client participation modeling used in the proof of this paper did not consider the Markov process, i.e., the client participation probabilities depend on the participation status in the last round. We believe that FOCUS still converges exactly under this scenario since the stochastic matrix modeling and push-pull strategy still hold for any realizations. However, the extension of the proof is non-trivial due to the correlation between stochastic matrices. We leave this and a more general arbitrary participation scenario for future research directions.
Future Works. The framework of stochastic matrices and time-varying graphs provides a novel tool for modeling arbitrary client participation and local update dynamics in FL. By leveraging this approach, we establish a formal connection between FL and the rich field of decentralized optimization. While this paper focused on FedAvg and the Push-Pull algorithm as initial examples, a promising avenue for future research is to adapt other sophisticated decentralized algorithms.
Acknowledgement
The authors gratefully acknowledge Edward Duc Hien Nguyen and Xin Jiang for the discussion that inspired the connection between time-varying graphs and client sampling. Research reported in this publication was supported by the National Institute Of General Medical Sciences of the National Institutes of Health under Award Number R16GM159671. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
@@ Appendix A Review of Decentralized Algorithms
It is well-known that the DGD with a fixed step size $\eta$ only converges to an ${\mathcal{O}}(\eta)$-sized neighborhood of the solution of the original algorithm (Yuan et al., 2016). This resembles the FedAvg algorithm exactly. Subsequently, several extract decentralized optimization algorithms have been proposed, such as Extra (Shi et al., 2015), exact-diffusion/NIDS (Yuan et al., 2018; Li et al., 2019), DIGing/Gradient tracking (Nedic et al., 2017), etc. The key advancement of these algorithms is their capability to achieve extra convergence under a fixed step size. They formulate the original sum-of-cost problem into a constrained optimization problem and then apply the primal-dual style approach to solve the constrained problem (Ryu and Yin, 2022).
While it is common in the analysis of decentralized algorithms to assume a static and strongly connected underlying graph structure, a significant body of research also investigates time-varying graph topologies (Lan et al., 2020; Saadatniaki et al., 2020; Assran et al., 2019; Ying et al., 2021; Nguyen et al., 2025). These studies often adopt one of two common assumptions regarding the dynamics of such graphs: either the union of graphs over any consecutive $\tau$ iterations or the expected graph is strongly connected (Nedić and Olshevsky, 2014; Koloskova et al., 2020). Whereas previous research on time-varying topologies often relies on specific graph assumptions, this paper takes a different approach. We model client sampling and local updates using a graph representation, thereby avoiding any presuppositions about the underlying graph structure.
Lastly, we want to point out that this work focuses quite differently from the decentralized FL work (Beltrán et al., 2023; Shi et al., 2023; Fang et al., 2024), which is more closely related to decentralized algorithms instead of FL settings.
@@ Appendix B Conventions and Notations
In this paper, $\|\cdot\|$ denotes (induced) $\ell_{2}$ norm for both vector and matrix usage while $\|\cdot\|_{F}$ denotes the Frobenius norm.
Appendix C Brief Review about Stochastic Matrices and their Properties
Before applying the formalism of stochastic matrices to FL algorithm, we review the key properties of row- and column-stochastic matrices, as a clear understanding of these concepts is essential for the analysis that follows. There are three different types of stochastic matrices.
  1. 1.
    A row stochastic matrix, also called right stochastic matrix, is a square matrix of nonnegative real numbers denoted as $R\in\mathbb{R}_{\geq 0}^{n\times n}$, with each row summing to 1, i.e., $R\mathds{1}=\mathds{1}$, where $\mathds{1}$ denotes the all-ones vector of size $n$.
  2. 2.
    A column stochastic matrix, also called left stochastic matrix, is a square matrix of nonnegative real numbers denoted as $C\in\mathbb{R}_{\geq 0}^{n\times n}$, with each column summing to 1, i.e., $\mathds{1}^{\mathsf{T}}C=\mathds{1}^{\mathsf{T}}$.
  3. 3.
    A doubly stochastic matrix is a square matrix $W\in\mathbb{R}_{\geq 0}^{n\times n}$ that is both row stochastic and column stochastic, i.e., $P\mathds{1}=\mathds{1}$ and $\mathds{1}^{\mathsf{T}}W=\mathds{1}^{\mathsf{T}}$.
Consensus property of Row Stochastic matrix
A row-stochastic matrix exhibits a consensus property. By the definition of row stochastic matrix $R$, we know $R\mathds{1}=\mathds{1}$. If the matrix $R$ is also primitive, the Perron-Frobenius theorem guarantees that all its other eigenvalues have a magnitude strictly less than 1. Consequently, the recursion
${\boldsymbol{x}}_{k+1}=R{\boldsymbol{x}}_{k},\;\;\mbox{where ${\boldsymbol{x}}_{0}\in\mathbb{R}^{N\times d}$}.$ (32)
converges to a consensus vector, ${\boldsymbol{x}}_{\infty}:=\lim_{k\to\infty}{\boldsymbol{x}}_{k}$, where all elements of $\boldsymbol{x}$$_{\infty}$ are identical. However, unlike the doubly stochastic matrix case, ${\boldsymbol{x}}_{\infty}\neq\frac{1}{n}\mathds{1}\mathds{1}^{\mathsf{T}}{\boldsymbol{x}}_{0}$ in general. Suppose that the corresponding left eigenvector of $R$ with eigenvalue is $p$, that is. $p^{\mathsf{T}}R=p^{\mathsf{T}}$. Then, it is easy to show that
${\boldsymbol{x}}_{\infty}=\lim_{k\to\infty}R^{k}{\boldsymbol{x}}_{0}=p\mathds{1}^{\mathsf{T}}{\boldsymbol{x}}_{0}$ (33)
In general, the Perron vector $p$ is not a uniform vector $\frac{1}{n}\mathds{1}$. Hence, the recursion of a row stochastic matrix typically converges to the consensus value of some weighted average of initial values.
To provide some concrete examples, consider the following row stochastic matrix:
$R=\frac{1}{2}\begin{bmatrix}1&0&1&0\\ 1&1&0&0\\ 1&0&1&0\\ 1&0&0&1\end{bmatrix},\;\;R^{\infty}=\frac{1}{2}\begin{bmatrix}1&0&1&0\\ 1&0&1&0\\ 1&0&1&0\\ 1&0&1&0\end{bmatrix}$ (34)
That is $p=[0.5,0,0.5,0]$ in this case.
Mass preservation property of Column Stochastic matrix
A column-stochastic matrix exhibits properties markedly different from a row-stochastic matrix. While repeatedly multiplying a vector by a row-stochastic matrix leads to consensus, this is generally not true for a column-stochastic matrix.
Instead, column-stochastic matrices possess a crucial property we call mass (or information) preservation. To understand this property, consider the following recursion:
${\boldsymbol{x}}_{k+1}=C_{k}{\boldsymbol{x}}_{k},\;\;\mbox{where ${\boldsymbol{x}}_{0}\in\mathbb{R}^{N\times d}$}.$ (35)
Note the subscript in $C_{k}$ indicates that a different column-stochastic matrix can be applied at each step $k$ of the recursion. If we imagine that each element ${\boldsymbol{x}}_{0}(i)$ represents the mass of the $i$-th object, then the recursion ${\boldsymbol{x}}_{k+1}=C_{k}{\boldsymbol{x}}_{k}$ can be interpreted as a mass redistribution process where the total mass is conserved at every step. To see that, left multiplying the $\mathds{1}^{\mathsf{T}}$ to both sides of above equation we get
$\mathds{1}^{\mathsf{T}}x^{k+1}=\mathds{1}^{\mathsf{T}}C_{k}x^{k}=\mathds{1}^{\mathsf{T}}x^{k}$ (36)
By induction, we can conclude $\mathds{1}^{\mathsf{T}}x^{k}=\mathds{1}^{\mathsf{T}}x^{0},\forall k$.
Finally, since a doubly-stochastic matrix is both row- and column-stochastic, it automatically inherits both of the properties discussed above. Combining the consensus-driving property (from being row-stochastic) with the mass-preservation property (from being column-stochastic) allows us to recover (unbiased) average consensus. The recursion converges to an average consensus, where every element of the final vector is equal to the average of the elements in the initial vector.
@@ Appendix C Brief Review about Stochastic Matrices and their Properties
Appendix D Proof of the Convergence of FedAvg under Arbitrary Client Participation
@@
This section presents a convergence analysis of the FedAvg algorithm with an arbitrary sampling/participation scheme. We focus on the strongly-convex case with a constant step size since this setting best illustrates the impact of client drift induced by local updates and bias introduced by non-uniform sampling. The following proof draws inspiration from and synthesizes several existing works (Wang and Ji, 2022; Koloskova et al., 2020; Li et al., 2020), adapting their insights to a decentralized framework. Leveraging this framework, we are able to present a more concise proof and provide a clearer conclusion. Unlike FOCUS, most FL algorithms require an extra bounded heterogeneity assumptions:
@@
[…]
@@
D.1 Reformation and Mixing Matrices
@@
See the definition of $R(S_{r})$ and $V(S_{r})$ in main context and example in Figure 6. A noteworthy observation from (Li et al., 2020) is that FedAvg can be equivalently reformulated without altering the trajectory of the server’s model. This reformulation considers activating all devices, where each device pulls the model from the server and performs a local update, while maintaining the same set of contributing clients for averaging. See Figure 7 as an illustration. Swap the order of $x$ and $y$ update, we arrive at